[Brahms-l] Shift report 20040222 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Sun Feb 22 2004 - 23:53:39 EST
------------------------------------------------------------
Date: Sun Feb 22 23:53:39 2004, Shift: 16:00-24:00
Supervisor: BHS, with: PJ
Log book#15 pages 65 - 65

Shift Summary:
------------------------------------------------------------
1. Good beam. There was 1-2 hours where they had problems storing.

2.
        FS 3deg 1/2A     MRS 40deg 1/2B
10497   180k             140k
10498   170k             185k
10501   122k             120k
10523   265k             235k
10524   Still Runnung

TOTAL   737k             680k

3.
the DAQ started acting up, but it was fixed by reseting the TCP2 and MTCP1 reciever board.
When I started my ped run (18:05), I discovered that pulser1 was already on. It's has probably been on since last ped run, which means that runs 10492, 104934, 10495, 10496, 10497, 10498 have been affected. We are not sure of the impact of this, so we havn't changed the quality of the data (if at all needed), we just noted it here (and in the elog).

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Feb 22 23:53:56 2004

This archive was generated by hypermail 2.1.8 : Sun Feb 22 2004 - 23:54:15 EST